(2+2y^2)dx+3xy^2dy=0


1
Expert's answer
2020-08-06T14:19:12-0400

Answer:

"\dfrac {3}{2y}ln|1+y^2| + ln|\dfrac {A}{x^2}| - 3y = 0"


Solution

"(2+2y^2)dx + 3xy^2dy = 0"

Separating terms gives:

"3xy^2dy = -2(1+y^2)dx"

dividing both sides by "x(1+y^2)" gives:

"\dfrac {3xy^2}{x(1+y^2)}dy = \dfrac {-2(1+y^2)}{x(1+y^2)}dx"


"=> \dfrac {3y^2}{1+y^2}dy = \dfrac {-2}{x}dx"


Reducing "\dfrac {y^2}{1+y^2}" will give:

"\dfrac {y^2}{1+y^2} = 1 - \dfrac{1}{1+y^2}"


"=> 3(1-\dfrac {1}{1+y^2})dy = \dfrac{-2}{x}dx"

Integrating both sides gives:

"\intop3(1-\dfrac {1}{1+y^2})dy = \intop\dfrac{-2}{x}dx"

"=> 3\intop(1-\dfrac {1}{1+y^2})dy = -2\intop\dfrac{1}{x}dx"

"=> 3(y-\dfrac{1}{2y}ln|1+y^2|=-2ln|x|+ln(A)"

"=> 3y - \dfrac{3}{2y}ln|1+y^2| = ln|x^{-2}| + ln(A)"

"=> 3y = \dfrac {3}{2y}ln|1+y^2| + ln|\dfrac {A}{x^2}|"


"\dfrac {3}{2y}ln|1+y^2| + ln|\dfrac {A}{x^2}| - 3y = 0"